# MultiConIR: Towards multi-condition Information Retrieval

### v2

@@ 3.3 Evaluation Metrics
Queries range from query1 to query10, each progressively incorporating 1 to 10 conditions. The candidate set comprises a Positive document that fully satisfies all conditions and a HN1 document, which is derived from the positive by modifying a single condition.
Complexity robustness is measured using Win Rate (WR) under various $k$, defined as:
$\text{WR}_{k}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}(S(q_{k},d^{+})>S(q_{k},d_{k-1})),$ (1)
@@
We evaluate performance using Win Rate between adjacent hard negatives:
$\begin{split}\text{WR}_{k,k-1}&=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\big{(}S(q_{10},d_{k})\\ &\quad>S(q_{10},d_{k-1})\big{)},\end{split}$ (2)
@@
$\begin{split}\text{FR}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\big{(}\mathbf{1}(S_{\text{inst}}(q_{10},d_{k})>S_{\text{inst}}(q_{10},d_{k-1}))\\ \neq\mathbf{1}(S_{\text{desc}}(q_{10},d_{k})>S_{\text{desc}}(q_{10},d_{k-1}))\big{)},\end{split}$ (3)
@@
Figure 3: Effect of Condition Quantity on Average Win Rate (reflecting complexity robustness). Win Rate measures Positive vs. Hard Negative success as query conditions increase. Blue, green and red lines represent Dense Retrievers, Rerankers and Red line respectively.
Figure 4: Relevance Monotonicity Distinction. Win Rate reflects the success rate between documents satisfying different numbers of conditions under a multi-condition query (query10), i.e., $d_{k}$ vs. $d_{k-1}$.
@@ 4 Experiments
We evaluate 13 representative retrieval models from diverse architectures and varying model sizes, including one spare retrieval model: BM25 (Robertson and Zaragoza, 2009); Two BERT-based retrieval models: gte-large-en-v1.5 (Li et al., 2023b) and jina-embeddings-v3 (Sturua et al., 2024); Seven LLM-based retrieval models: NV-Embed-v2 (Lee et al., 2024), bge-en-icl (Li et al., 2024), gte-Qwen2-7B-instruct (Li et al., 2023b), gte-Qwen2-1.5B-instruct (Li et al., 2023b), e5-mistral-7b-instruct (Wang et al., 2024a), GritLM-7B (Muennighoff et al., 2024), LLM2Vec (BehnamGhader et al., 2024); Three reranking models: bge-reranker-v2-m3 (Chen et al., 2024), bge-reranker-v2-gemma (Chen et al., 2024), FollowIR-7B (Weller et al., 2024a). Details of each model are provided in Appendix A.
@@
jina-embeddings-v3 76.09 71.26 71.84 71.04 65.59 65.65 64.75 64.24 64.62 60.71 15.38
gte-large-en-v1.5 75.87 77.26 73.79 70.22 70.71 67.40 67.53 64.97 65.36 61.36 14.51
NV-Embed-v2 80.53 80.32 78.81 75.70 75.68 72.61 73.28 71.54 70.00 68.02 12.51
bge-en-icl 83.42 80.65 78.44 76.77 74.54 73.00 74.23 73.25 69.70 68.00 15.42
gte-Qwen2-7B-instruct 70.75 72.22 69.99 68.51 65.20 63.53 62.22 62.20 59.15 56.17 14.58
gte-Qwen2-1.5B-instruct 73.64 74.97 72.23 71.37 69.94 67.46 66.92 64.64 63.65 58.68 14.96
e5-mistral-7b-instruct 75.05 70.85 68.18 67.45 63.70 61.60 59.70 59.07 57.85 58.12 16.93
GritLM-7B 82.08 80.32 78.38 76.40 76.40 73.50 75.69 74.62 74.53 75.95 6.13
LLM2Vec 83.13 77.42 75.49 75.48 72.49 72.56 70.56 70.73 68.71 67.00 16.13
Reranker
bge-reranker-v2-m3 87.14 85.56 78.62 76.05 74.29 68.41 67.86 59.48 55.59 44.87 42.27
bge-reranker-v2-gemma 91.07 90.02 86.70 84.99 83.17 79.00 75.89 72.29 67.11 56.09 34.98
followIR 83.41 79.72 76.25 74.60 70.12 67.94 62.62 55.93 48.59 43.52 39.89
@@ 4.1 Results for Complexity Robustness
Table 1 presents the average Win Rate scores for evaluating complexity robustness across five datasets [^3] Individual dataset performances are detailed in the Appendix D.1.. To visualize how model performance evolves as the number of query conditions increases, we plot a line graph with Gaussian smoothing (window size = 1), as shown in Fig. 4. The results reveal several notable trends:
Performance decline with increasing query conditions
As the number of conditions in the query increases, the performance of both retrieval and reranking models declines. This suggests that with more conditions, models struggle to accurately distinguish between positives and HNs. Among all models, GritLM-7B exhibits the lowest performance degradation, declining by only 6.13% from Query1 to Query10. NV-Embed-v2 follows with a relatively low decline of 12.51%.
Reranking models exhibit steeper performance drop
Initially, reranking models outperform retrieval models with single-condition queries. However, as the number of conditions increases, their performance declines more sharply. As shown in Table 1, for Query1, rerankers almost outperformed all retrieval models. bge-reranker-v2-gemma achieves a Win Rate of 91.07% at Query1 but drops to 56.09% at Query10, a drastic 34.98% decrease. Similarly, bge-reranker-v2-m3 declines from 87.14% to 44.87%, a drop of 42.27%, and followIR shows a drop from 83.41% to 43.52%, with a drop of 39.89%. In contrast, retrievers exhibit an average Win Rate decline of 14.06%.
Sparse retrieval models show performance improvement
Unlike dense retrievers and rerankers, BM25’s performance improves with query complexity, increasing from 28.59% (Query1) to 39.87% (Query10). This is likely due to its lexical matching mechanism, where more query conditions lead to greater lexical overlap with candidate documents, enhancing retrieval effectiveness. However, BM25’s overall win rate remains lower than dense retrievers, and positives rank below HNs (win rate < 0.5), suggesting that hard negatives consistently outrank positives. This may stem from dataset construction, where queries and HNs were generated or paraphrased by GPT-4o, potentially introducing lexical bias [^4] Appendix E further explores challenges in using generated datasets..
@@ 4.2 Results for Relevance Monotonicity
Fig. 4 illustrates the trend of average Win Rate in the multi-condition retrieval setting of Task 2555The complete results are provided in Appendix D.2, which evaluates the model’s ability to distinguish the relevance hierarchy among documents with varying conditions. Several key observations can be made:
Relevance monotonicity failure
As documents become increasingly hard (i.e., satisfying more conditions in the query), retrieval and reranking models struggle to distinguish between $d_{k}$ and $d_{k-1}$, leading to a decline in Win Rate performance. This failure emphasizes the challenge of preserving relevance monotonicity in multi-condition retrieval settings and highlights a gap in current model capabilities when handling complex queries.
Sensitive to “exact match” and “complete mismatches”
@@ 4.3 Results for Format Invariance
Table 2 presents the Flip Rate induced by query format variations. Most models exceed 10%, indicating a substantial impact of query formulation on retrieval performance. BM25 records a 16.19% Flip Rate, which can serve as a baseline for the lexical differences caused by query format variations.
Dense retrieval models show relatively lower sensitivity, with Flip Rates between 8% to 16%. GritLM-7B (8.21%), NV-Embed-v2 (9.12%), and LLM2Vec (9.78%) exhibit less variation, suggesting higher robustness to query styles.
In contrast, reranking models show significantly higher sensitivity to query format changes, with Flip Rates exceeding 20%. The highest Flip Rate observed is 33.81% for bge-reranker-v2-m3, confirming their susceptibility to query formulation changes.
@@
(b) Retrieval Performance when padding to 512 words
Figure 5: Retrieval performance when padding the document set to 512 words and 1024 words.
@@ 5 Analysis
Our experiments reveal notable differences between retrieval and reranking models across the three tasks. Specifically, retrieval models demonstrate greater robustness to query complexity, better preservation of relevance monotonicity, and stronger invariance to query format changes compared to rerankers. We analyze these differences through the lens of relevance computation and attention mechanisms.
Retrieval models: robustness induced by bidirectional attention and dual-encoder.
Retrieval models typically employ a dual-encoder architecture, where queries and documents are independently encoded before computing their similarity using dot-product or cosine similarity. This independent computation ensures that the generation of query and document embeddings remains unaffected by each other. At the same time, bidirectional attention enables the model to better capture the overall semantic meaning of the query.
@@
Another critical factor is the training datasets. Many dense retrieval models are trained on a mixture of retrieval-specific and general textual datasets (Lee et al., 2024; BehnamGhader et al., 2024; Wang et al., 2024a), which enhances their generalization across diverse retrieval scenarios and query styles. This, in turn, improves their robustness against query complexity.
Reranking models: sensitivity to query complexity.
Figure 6: Attention heatmap of cross-encoder model (bge-reranker-m3): effect of increasing query conditions. When query conditions are few (1-3), the cross-encoder effectively attends to relevant conditions in the document, as the number of conditions increases, attention becomes more diluted.
Reranking models compute relevance by jointly processing the query and document, primarily through: (1) cross-encoders, which perform token-level relevance comparison through cross-attention (e.g., bge-reranker-v2-m3), and (2) generative models that estimate relevance using LLM-based agents (e.g., bge-reranker-v2-gemma and FollowIR). Both architectures rely on deep query-document interaction, making them more sensitive to input complexity, such as changes in condition quantity and query format.
Our experiments show that rerankers outperform retrieval models in ranking effectiveness for single-condition queries, suggesting their advantage in capturing fine-grained query-document relevance for short and simple queries. However, as query complexity increases, their performance declines more sharply, eventually falling behind retrieval models. Table 1 shows that rerankers’ average Win Rate drops by 39.05%, more than twice the decline of retrieval models (14.06%).
A case study in Fig. 6 provides further insight into this decline by analyzing the attention heatmap of a cross-encoder reranker. When the query contains only a few conditions (e.g., 1–3), the model effectively aligns each condition with relevant document sentences. However, as conditions increase, attention becomes unevenly distributed, with earlier conditions receiving more focus while later ones are overlooked. This imbalance likely contributes to the sharp drop in Win Rate observed in Task 2—as later conditions receive insufficient attention, the model struggles to distinguish relevance variations in corresponding document sentences.
These findings highlight a key limitation in existing reranking models: their reliance on fine-grained query-document interactions becomes a bottleneck in multi-condition scenarios. While reranking models outperform retrieval models for queries with fewer conditions, their effectiveness diminishes as query complexity grows.
5.2 The Effect of Document Side
Previous experiments focused on query-side variations. We now investigate how changes on the document side, particularly document length, affect retrieval effectiveness.
Figure 5 shows the impact of document length (padded to 512 and 1024 words) on relevance monotonicity [^6] The complete results and the padding filler are provided in Appendix D.3. Compared to original documents (Fig.4), increasing document length does not change the overall trend in Task 2—as hard negatives (HNs) become more challenging, models struggle more to distinguish relevance between documents satisfying $k$ vs. $k-1$ conditions. This decline is more pronounced when document length increases to 1024 words.
Rerankers are particularly affected, showing a steeper win rate drop than retrieval models. Since they process queries and documents jointly, longer inputs likely increase attention dispersion in cross-encoders and complexity in generative models, weakening relevance estimation.
Among retrieval models, GritLM and Nv-embed demonstrate strong robustness to document length variations. At 512 words, GritLM (66.66%) and Nv-embed (62.91%) achieve the highest average win rate. When document length is increased to 1024 words, Nv-embed (63.41%) outperforms GritLM (59.46%), indicating its ability to retain essential information in longer documents. This suggests that Nv-embed is more effective at capturing key content even when document length increases.
5.3 Robustness of GritLM and NV-Embed
@@
GritLM and hybrid attention.
GritLM exhibits strong robustness in multi-condition retrieval, which we hypothesize is due to its hybrid attention mechanism, integrating causal attention for generative task and bidirectional attention for embedding learning.
Recent work has incorporated bidirectional attention into LLM-based retrievers, such as LLM2Vec (BehnamGhader et al., 2024) and Nv-Embed (Lee et al., 2024), which discard causal attention after converting LLMs into embedding models. In contrast, GritLM preserves causal attention, maintaining generative reasoning capabilities alongside representation learning. This likely explains its superior adaptability in multi-condition retrieval. Empirical results further validate this advantage: GritLM achieves state-of-the-art performance in complex instruction-following retrieval (Zhou et al., 2025) and ranks second in reasoning-intensive retrieval tasks (Su et al., 2024), reinforcing the importance of hybrid attention in embedding-based retrieval.
Nv-Embed and latent attention layer pooling.
Nv-Embed performs suboptimally in most multi-condition retrieval tasks, which we hypothesize is due to its Latent Attention Layer Pooling strategy. Common embedding methods include: 1) mean pooling, and 2) the last <EOS> token embedding. Mean pooling averages all token representations to generate a global representation, but potentially dilute critical information. Last <EOS> token embedding relies solely on the final token, which could introduce recency bias.
A case study in Fig. 7 examines how different pooling methods handle condition placement in queries. The results show that mean pooling leads to decreasing similarity scores as the condition moves further back, while <EOS> pooling exhibits an increasing trend, confirming its recency bias. In contrast, latent attention layer pooling maintains relatively stable performance across different condition positions, indicating improved information retention.
We validated the robustness of Nv-Embed in handling longer documents in Section 5.2. Notably, at 1024 words, it maintains consistent monotonic relevance, preserving key information more effectively than other models. These findings suggest that its latent attention layer pooling potentially mitigates information loss, enhancing retrieval stability in multi-condition scenarios.
@@ 6 Conclusion
In this paper, we introduce MultiConIR, the first benchmark specifically designed to evaluate retrieval models in multi-condition scenarios. Unlike existing datasets that primarily focus on single-condition queries from search engines, MultiConIR captures real-world complexity by incorporating five diverse domains: books, movies, people, medical cases, and legal documents.
We propose three tasks to systematically assess retrieval and reranking models on multi-condition robustness, monotonic relevance ranking, and query format sensitivity. Through extensive experiments, we analyze model performance across architectures and sizes, examining their ability to distinguish fine-grained conditions, preserve semantic information, and adapt to contextual variations.
Our findings reveal that existing retrieval and reranking models struggle with multi-condition retrieval, with rerankers suffering severe performance degradation as query complexity increases. We further investigate the performance gap between retrieval and reranking models, exploring potential reasons for these discrepancies. Finally, we highlight the strengths of GritLM and Nv-Embed, which demonstrate enhanced adaptability to multi-condition queries, offering insights for future retrieval models.

### v3

@@ 3.3 Evaluation Metrics
Conventional IR metrics—e.g., Precision@1, NDCG@kk—only confirm whether a highly relevant document appears early, but they cannot distinguish how well a model orders candidates that satisfy different numbers of query conditions. We therefore propose Win Rate: the proportion of pairwise comparisons in which a candidate that fulfils more conditions ranks above one that fulfils fewer. We further discuss the difference of Win Rate and traditional IR metrics in the Appendix D
@@
Queries range from query1 to query10, each progressively incorporating 1 to 10 conditions. The candidate set comprises a Positive document that fully satisfies all conditions and a HN1 document, which is derived from the positive by modifying a single condition. Complexity robustness is measured using Win Rate (WR)  [^3] In Complexity Robustness evaluation, $\text{WR}_{k}$ and Precision@1 are numerically equivalent. under various $k$, defined as:
$\text{WR}_{k}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}(S(q_{k},d^{+})>S(q_{k},d_{k-1})),$
@@
We evaluate performance using $\text{WR}_{k,k-1}$ between adjacent hard negatives:
$\begin{split}\text{WR}_{k,k-1}&=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\big{(}S(q_{10},d_{k})\\ &\quad>S(q_{10},d_{k-1})\big{)},\end{split}$
@@
$\begin{split}\text{FR}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\big{(}\mathbf{1}(S_{\text{inst}}(q_{10},d_{k})>S_{\text{inst}}(q_{10},d_{k-1}))\\ \neq\mathbf{1}(S_{\text{desc}}(q_{10},d_{k})>S_{\text{desc}}(q_{10},d_{k-1}))\big{)},\end{split}$
@@ 4 Experiments
We evaluate 15 representative retrieval and reranking models from diverse architectures and varying model sizes, including sparse retrieval model: BM25 (Robertson and Zaragoza, 2009); BERT-based retrieval models: gte-large-en-v1.5 (Li et al., 2023b) and jina-embeddings-v3 (Sturua et al., 2024); LLM-based retrieval models: NV-Embed-v2 (Lee et al., 2024), bge-en-icl (Li et al., 2024), gte-Qwen2-7B-instruct (Li et al., 2023b), gte-Qwen2-1.5B-instruct (Li et al., 2023b), e5-mistral-7b-instruct (Wang et al., 2024a), GritLM-7B (Muennighoff et al., 2024), LLM2Vec (BehnamGhader et al., 2024); Point-wise reranking models: bge-reranker-v2-m3 (Chen et al., 2024), bge-reranker-v2-gemma (Chen et al., 2024), FollowIR-7B (Weller et al., 2024a); Fine-tuned list-wise reranker: RankZephyr (Pradeep et al., 2023); Advanced LLM for zero-shot ranking: GPT-4o (OpenAI, 2024). Details of each model are provided in Appendix A.
@@
jina-embeddings-v3 76.09 71.26 71.84 71.04 65.59 65.65 64.75 64.24 64.62 60.71 $\downarrow$15.38
gte-large-en-v1.5 75.87 77.26 73.79 70.22 70.71 67.40 67.53 64.97 65.36 61.36 $\downarrow$14.51
NV-Embed-v2 80.53 80.32 78.81 75.70 75.68 72.61 73.28 71.54 70.00 68.02 $\downarrow$12.51
bge-en-icl 83.42 80.65 78.44 76.77 74.54 73.00 74.23 73.25 69.70 68.00 $\downarrow$15.42
gte-Qwen2-7B-instruct 70.75 72.22 69.99 68.51 65.20 63.53 62.22 62.20 59.15 56.17 $\downarrow$14.58
gte-Qwen2-1.5B-instruct 73.64 74.97 72.23 71.37 69.94 67.46 66.92 64.64 63.65 58.68 $\downarrow$14.96
e5-mistral-7b-instruct 75.05 70.85 68.18 67.45 63.70 61.60 59.70 59.07 57.85 58.12 $\downarrow$16.93
GritLM-7B 82.08 80.32 78.38 76.40 76.40 73.50 75.69 74.62 74.53 75.95 $\downarrow$6.13
LLM2Vec 83.13 77.42 75.49 75.48 72.49 72.56 70.56 70.73 68.71 67.00 $\downarrow$16.13
Fine-tuned Reranker
bge-reranker-v2-m3 87.14 85.56 78.62 76.05 74.29 68.41 67.86 59.48 55.59 44.87 $\downarrow$42.27
bge-reranker-v2-gemma 91.07 90.02 86.70 84.99 83.17 79.00 75.89 72.29 67.11 56.09 $\downarrow$34.98
followIR 83.41 79.72 76.25 74.60 70.12 67.94 62.62 55.93 48.59 43.52 $\downarrow$39.89
RankZephyr 92.72 90.29 88.38 87.69 84.57 80.88 78.93 75.99 72.39 66.84 $\downarrow$25.88
Zero-shot LLM for Ranking
GPT-4o 95.49 94.89 93.71 92.11 90.81 89.08 88.43 88.08 86.82 85.26 $\downarrow$9.23
@@ 4.1 Results for Complexity Robustness
Table 2 presents the average Win Rate scores for evaluating complexity robustness across five datasets. The results reveal several notable trends:
Performance decline with more conditions
As the number of conditions in the query increases, the performance of both retrieval and reranking models declines. This suggests that with more conditions, models struggle to accurately distinguish between positives and HNs. Among all models, GPT-4o maintained the highest win rate from Query1 to Query10, with its performance declining by 9.23%. GritLM-7B exhibits the lowest performance degradation of 6.13%.The remaining models all exceeded 10% decline.
Rerankers exhibit steeper performance drop
As shown in Table 2, fine-tuned rerankers outperform retrievers with single-condition queries. However, as the number of conditions increases, their performance declines more sharply. Eventually, rerankers even fell behind some retrievers. The Win Rates for all rerankers declined by over 25%, with an average decline of 35.76%. For retriever models, the average decline was 14.06%.
@@ 4.2 Results for Relevance Monotonicity
Fig. 4 illustrates the trend of average $\text{WR}_{k,k-1}$ in the multi-condition retrieval setting of Task 2, which evaluates the model’s ability to distinguish the relevance hierarchy among documents with varying conditions. The complete results are provided in Table 13. Several key observations can be made:
Relevance monotonicity struggle
As documents become increasingly hard (i.e., satisfying more conditions in the query), it becomes harder for retrieval and reranking models to accurately distinguish $d_{k}$ and $d_{k-1}$, leading to a decline in Win Rate performance. This failure emphasizes the challenge of preserving relevance monotonicity in multi-condition retrieval settings and highlights a gap in current model capabilities when handling complex queries.
Sensitive to exact match and complete mismatches
@@ 4.3 Results for Format Invariance
Table 3 presents the Flip Rate induced by query format variations. GPT-4o showed the lowest flip rate of 6.98%, showcasing the robustness of advanced LLMs against variations in query style. Additionally, most models exceed 10%, indicating a substantial impact of query formulation on retrieval performance. Table 3 presents the complete results of Format Invariance.
@@
Dense retrieval models show relatively lower sensitivity than rerankers, with Flip Rates between 8% to 16%. GritLM-7B (8.21%), NV-Embed-v2 (9.12%), and LLM2Vec (9.78%) exhibit less variation. In contrast, reranking models show significantly higher sensitivity to query format changes, with Flip Rates exceeding 20%. The highest Flip Rate observed is 33.81% for bge-reranker-v2-m3.
Figure 4: Relevance Monotonicity Distinction. Win Rate reflects the success rate between documents satisfying different numbers of conditions under a multi-condition query.
@@
(b) Retrieval Performance when padding to 1024 words
Figure 5: Retrieval performance when padding the document set to 512 words and 1024 words. Rerankers are highly sensitive to increases in document length, showing rapid performance degradation, whereas retrievers remain comparatively robust.
@@ 5 Analysis
Our experiments reveal notable differences between retrievers and rerankers across the three tasks. Fine-tuned rerankers exhibited excellent ranking performance under single-condition queries, but their efficacy rapidly diminished as the number of conditions increased. Retrievers demonstrate greater robustness under multi-condition queries and query style variations.
@@
We hypothesize that one contributing factor to these performance disparities lies in the training datasets. Many dense retrieval models are trained on a mixture of retrieval-specific and general textual datasets (Lee et al., 2024; BehnamGhader et al., 2024; Wang et al., 2024a). Such diverse training enhances their generalization across various retrieval scenarios and query styles, which, in turn, improves their robustness against query complexity.
Beyond training data, we posit that the distinct input processing mechanisms also contribute to the observed performance differences. Retriever models typically employ a bi-encoder architecture, processing queries and documents independently. Conversely, rerankers, which process a concatenation of the query and document as a single input, appear more susceptible to input complexification—whether arising from an increase in query conditions or changes in query style. To validate this speculation, we re-evaluated the Win Rate for the relevance monotonicity task under multi-condition queries using documents that padded to 512 and 1024 words. Results as shown in Fig.5 , revealed that fine-tuned rerankers are highly sensitive to such increases in document length, which further illustrates the sensitivity of rerankers to complex input. In contrast, retriever models demonstrated greater robustness to length modifications.
5.2 Condition Position Impact on Focus
@@
Our experimental findings indicate that the position of a condition significantly influences the model’s subsequent similarity judgment. This phenomenon was observed consistently across both retriever and reranker models.
To illustrate position effects in retriever models, we conducted a targeted study on how relocating a single condition within the query influences its similarity score. We selected representative retriever models employing distinct pooling strategies: mean pooling, <EOS> pooling, and latent layer pooling. The results, as depicted in Fig.6, revealed that models utilizing mean pooling tend to weight the early tokens most heavily: similarity drops steadily as the condition is shifted toward the tail of the query. <EOS> shows the opposite bias, emphasising the final tokens. Latent layer pooling heightened focus on both the beginning and end of the query, with comparatively less focus on the middle.
Figure 7: Attention heatmap of cross-encoder model (bge-reranker-m3).
Similarly, for reranker models, we selected a cross-encoder model (bge-reranker-m3) to visualize the attention heatmap. Fig.7 shows a non-uniform distribution of attention across different token positions. This implies that these rerankers tend to assign differential focus to specific conditions or tokens within the concatenated query-document input, rather than distributing their attention uniformly across all elements.
@@ 6 Conclusion
In this work, we introduced MultiConIR, a novel benchmark designed to rigorously evaluate information retrieval models in realistic multi-condition scenarios, a critical area where existing evaluation frameworks are lacking. Through three specifically designed tasks—complexity robustness, relevance monotonicity, and query format sensitivity—conducted across five diverse domains. Experiments revealed that existing models struggle with multi-condition retrieval, with their performance degrading as the number of conditions increases; rerankers excel for single-condition queries but fail in multi-condition scenarios. Notably, rerankers are more sensitive to complex inputs. GPT-4o outperforms specialised IR systems, exposing a performance gap in handling complex information needs.
Our findings highlight an urgent need for new modeling approaches and training paradigms specifically tailored for robust multi-condition understanding. MultiConIR serves as a valuable resource to drive this research, benchmark progress, and ultimately propel information retrieval systems towards a more sophisticated, human-like comprehension of complex information needs.